OpenStudy (anonymous):

cos^2 theta (1+tan^2 theta)

OpenStudy (callisto):

Recall: 1 + tan^2 theta = sec^2 theta sec^2 theta = 1/ (cos^2 theta)

OpenStudy (anonymous):

would the answer be 1

OpenStudy (callisto):

It should be~

OpenStudy (anonymous):

1-2sin^2theta

OpenStudy (anteater):

1 + tan^2 theta = sec^2 theta sec^2 theta = 1/cos^2 theta So, cos^2 theta (1 + tan^2 theta) = cos^2 theta/cos^2 theta = 1